OpenStudy (anonymous):

Solve for x. |2x + 2| = 14 Can someone please show me how to do absolute value equations?

Parth (parthkohli):

\( \color{Black}{\Rightarrow |2x + 2| = 14 \Longrightarrow 2x + 2 = \pm 14 }\)

Parth (parthkohli):

So we have two equations \( \color{Black}{\Rightarrow 2x + 2 = 14 }\) AND \( \color{Black}{\Rightarrow 2x + 2 = -14}\) solve both and you get both solutions

OpenStudy (anonymous):

So, I need to do the equation but write it down twice with the positive and negative sign?

Parth (parthkohli):

Yep

OpenStudy (anonymous):

Thank you